OpenStudy (anonymous):

The course for a walkathon is 6 miles long. If Chloe wants to walk 50 miles during the walkathon, how many times must she walk the course?

OpenStudy (anonymous):

8.3333333

OpenStudy (anonymous):

6x = 50 x = 50/6 x = 8(1/3) or 8.3333333333 So 9 full times to walk more than 50 miles The answer is 9 =)

OpenStudy (anonymous):

Can you walk 1/3 of a marathon? xD It depends I guess both are right. Technically speaking it's 8.333333333 times; but generally I've always found they're looking for that rounded up which would be 9 times - But I'm guessing either will be correct

OpenStudy (anonymous):

i dont walk 1/3 of a marathon...i run haha

OpenStudy (anonymous):

Oh see I'm lazy... I would bus it =)

OpenStudy (anonymous):

i like we are using her page to talk lol

OpenStudy (anonymous):

Haha the question is answered ^.^ Always saying not to waste paper... Technically not paper but it put my mind to ease =)